WHat is the value you of nCr when n= r+1?
Answer by jim_thompson5910(35256)

n C r = (n!)/(r!(n-r)!)


n C r = ((r+1)!)/(r!(r+1-r)!)


n C r = ((r+1)!)/(r!(1)!)


n C r = ((r+1)!)/(r!*1)


n C r = ((r+1)!)/(r!)


n C r = ((r+1)*r!)/(r!)


n C r = r+1


So, n C r = r+1 when n = r+1